OpenStudy (anonymous):

Can I have someone's opinion Its a very easy question. I just want to know if i got everything down.

OpenStudy (anonymous):

@wolfe8

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

Are the two figures below similar? Explain why or why not using complete sentences. Provide evidence to support your claims.

OpenStudy (anonymous):

Hello, using the same answer as the last one. This time we can use SSS postulate. "SSS in same proportion (side side side) All three pairs of corresponding sides are in the same proportion" So in order to prove, you must calculate the length of each side For example, I calculate AF = |-4 - (-1) | = 3, GL = | 2 - 4 | = 2. Then you have the base ratio AF/GL = 3/2. B(-3,0), A(-4,-2) ==> \[AB = \sqrt{(-4-(-3))^2+(0-2)^2} = \sqrt{1^2+2^2} = \sqrt{5}\] Similarly, \[HG = \sqrt{2}\] Then ratio AB/HG = \[\frac{ \sqrt{5} }{ \sqrt{2} } = \frac{ 5 }{ 2 }\] You need all of side is at the same ratio = 5/2 so you must calculate the rest one. Good luck.

OpenStudy (anonymous):

@wolfe8

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

@someoneelse

OpenStudy (anonymous):

@chenping

OpenStudy (chenping):

Haha, well. I'm not sure as I have left Modern Math so long ago.. :P All I remembered is that transformation. Try joining all 8 points corresponding, you will find an origin point for the transformation (damn, I 4got the term..). So, I think that they are similar but also different :P Confused right? They are similar because the left 1 is actually just a transformation of the right 1. Yet, they are different in terms of scale. It's x2 scale. Haha, hope this may help :P

OpenStudy (wolfe8):

I'll look for a similar question I helped with a few days ago

OpenStudy (wolfe8):

http://openstudy.com/users/adrynicoleb#/updates/52b39b73e4b03e81ff9596b5 Hope that helps.

OpenStudy (agent0smith):

Look at the angles - do the angles look like they're all exactly the same size in both figures?

OpenStudy (agent0smith):

For example: Is angle JKL exactly the same size as angle DEF?

OpenStudy (anonymous):

one reason will be enough for proving that they are not similar, look at distances\[\frac{CD}{IJ} \neq \frac{DF}{JL}\]

OpenStudy (anonymous):

another reasoning comes from what agent0smith mentioned above